### v1

# Finslerian structure of black holes

Abstract
Recently, there has been an increasing interest in the Finslerian interpretation of null geodesics in the exterior regions of stationary black holes, particularly through the Zermelo navigation problem and the Randers metric. In this work, we show that recent mathematical advancements in the critical and strong Zermelo navigation problems and their connections to Kropina and Lorentz-Finsler metrics enable the extension of this Finslerian framework to encompass horizons and their interior regions. The Finslerian indicatrix, a key element of this framework, serves as an effective tool for identifying frame-dragging effects and the location of horizons and ergosurfaces. We illustrate our results with explicit physical examples, focusing on spherically symmetric and Kerr black holes. Our findings provide new insights into the “river model” of black holes, offering enhanced visual representations of null geodesics on the ergosurfaces, horizons, and within their interior regions.
I Introduction
Finsler geometry is a classical and well-established branch of mathematics with a vast body of literature. Recently, it has experienced significant developments and has found a wide range of applications in Physics and other applied sciences. While its main foundational ideas were already present in Riemann’s Habilitationsvortrag of 1854, its origins are typically traced back to Paul Finsler’s pioneering dissertation [1]. A modern perspective on the topic can be found, for instance, in [2]. For an abbreviated historical account, see [3]. Roughly, Finsler geometry is a Riemannian geometry in which the element of arc length $ds=F(\mathbf{x},d\mathbf{x})$ is not subjected to the usual quadratic restriction $F^{2}(\mathbf{x},d\mathbf{x})=a_{ij}dx^{i}dx^{j}$, with $a_{ij}$ standing for the ordinary Riemannian metric tensor. This restricted case has been historically referred to simply as Riemannian geometry, and any other form of $F(\mathbf{x},d\mathbf{x})$ is now termed Finslerian. Numerous Riemannian geometrical notions, ideas, and quantities have corresponding counterparts in Finsler geometry, with virtually every problem in Riemannian geometry having an equivalent formulation in the Finslerian framework.
Arguably, the simplest explicit example of a Finsler geometry in Physics is the Randers spacetime [4] with its “eccentric metric” now known as the Randers metric
$ds=F(\mathbf{x},d\mathbf{x})=\sqrt{a_{\mu\nu}(x^{\sigma})dx^{\mu}dx^{\nu}}+b_{\mu}(x^{\sigma})dx^{\mu},$ (1)
where $\mathbf{x}$ is a point with coordinates $(x^{\sigma})$ in the spacetime manifold $\mathcal{M}$ and $d\mathbf{x}\in T_{\mathbf{x}}\mathcal{M}$. We denote the components of $d\mathbf{x}$ by $(dx^{\mu})$, following the standard physical convention in which Greek indices run over all $n+1$ spacetime coordinates, while Latin indices are limited to the $n$ spatial dimensions. In the mathematical literature, $F(\mathbf{x},d\mathbf{x})$ is typically referred to as a (Minkowski) norm rather than a metric. In this work, however, we refer to it as a metric, following Randers’ original terminology and ensuring consistency with the conventions commonly adopted in the physical literature. In the Randers metric (1), $a_{\mu\nu}$ is a standard metric tensor (Lorentzian, in the case of Spacetime Physics), and $b_{\mu}$ is a one-form (covariant vector) that encodes a directional dependence of the geometry. If $b_{\mu}=0$, the underlying geometry reduces to the standard Riemannian/Lorentzian one associated with the metric tensor $a_{\mu\nu}$. Notice that $F(\mathbf{x},d\mathbf{x})=F(\mathbf{x},-d\mathbf{x})$ for generic vectors $d\mathbf{x}\in T_{\mathbf{x}}\mathcal{M}$ if, and only if, $b_{\mu}=0$. In other words, a Finsler metric is symmetric if and only if it is Riemannian. Notwithstanding, generic Finsler metrics are positively homogeneous of degree one, i.e., $F(\mathbf{x},\lambda d\mathbf{x})=\lambda F(\mathbf{x},d\mathbf{x})$ for all $\lambda\geq 0$. All Finslerian-geometric objects derived from (1), such as the fundamental form, connections, covariant derivative, and curvature, are defined starting from $F^{2}(\mathbf{x},d\mathbf{x})$ analogously to the Riemannian case (see [2] and references therein for further details). Positive definite metrics of the Randers type (1) arise naturally [5] in the classical optimization problem known as the Zermelo navigation problem [6], which consists in finding minimal (non-relativistic) time trajectories in curved spaces under the influence of vector drifts (winds). Many recent advancements in Finsler geometry and its applications in Physics are closely related to the Zermelo problem and its solutions involving metrics such as those of Randers type. For a comprehensive review on the the recent mathematical literature, see [7]. This work will further explore the different types of Zermelo navigation problems and their connections to other Finslerian metrics in Spacetime Physics.
From a physical standpoint, one does not expect distinguished local spatial directions. Hence, one would expect the emergence of Finslerian metrics like (1) in Spacetime Physics only when treating problems where the temporal direction is explicitly involved. In fact, this was the main motivation behind Randers’ introduction of a spacetime metric with asymmetrical properties of the type (1) in the early forties [4]. Moreover, the Principle of Equivalence imposes restrictions on possible forms of a Finsler metric from the Spacetime Physics perspective [8]. Interestingly, positive definite metrics of type (1) often appear in specific spacetime problems restricted to spatial sections, such as the description of causal structures. In particular, the propagation of light rays in stationary spacetimes can be effectively described by a Randers spatial metric. These issues were carefully and comprehensively considered for black holes in [9]. The extension for nonstationary but stably causal spacetimes admitting a globally defined time coordinate was considered in [10]. More recently, [11] has shown that the Zermelo problem can be used to generate spacetimes with specific causal properties. To date, only the propagation of light rays in the exterior region of black holes has been described in Finslerian terms in the Physics literature. Extending the Finsler structure across the event horizon is a physically relevant problem, as recognized in [11]. We highlight that recent applications of Finsler geometry in other problems of General and Special Relativity can be found, e.g., in [12, 13, 14, 15, 16, 17, 18, 19, 20].
In this paper, we demonstrate how recent mathematical progress [7, 21, 22, 23] can be explored to interpret light propagation within the interior region of black holes in Finslerian terms. The key ingredients are the Kropina and Lorentz-Finsler metrics, which appear in certain extended versions of the Zermelo navigation problem. We show, through explicit and physically relevant examples, that the Kropina metric adequately describes the causal properties of ergosurfaces in stationary black holes, while the Lorentz-Finsler metric captures, in the region where it is well defined, the causal structure of the interior (non-stationary) region of black holes. In this way, we show that it is indeed possible to have, based on recent mathematical developments, a complete Finslerian description of the causal structure of black holes, shedding new light on phenomena such as the one-directional membrane behavior of horizons [24] and the river model paradigm of black holes [25].
The paper is organized as follows. In the next section, we review recent mathematical developments related to the Kropina and Lorentz-Finsler metrics and compile the results that are instrumental in describing the causal structure of black holes, which is explored in Section III. The final section is dedicated to our concluding remarks, which include brief comments on Fermat’s principle and applications to analog spacetimes. We adopt $G=c=1$ units throughout this work.
II Finsler geometry and the Zermelo problem
We begin by revisiting the classical Zermelo navigation problem in curved space, as it provides the essential foundation for our approach in Section III. Consider a non-relativistic particle moving with constant (unit) speed in an $n$-dimensional Riemannian manifold $(\mathcal{M},h_{ij})$ subject to a time-independent vector field drift (wind) $W^{i}(x^{j})$ such that $||W||^{2}=h_{ij}W^{i}W^{j}<1$ (weak drift). The Zermelo problem consists in finding the trajectory that minimizes the transit time for a particle traveling between two distinct points of $\mathcal{M}$. This minimal time trajectory corresponds to the geodesic of a positive definite Randers metric (1) with [5]
$a_{ij}=\frac{1}{\lambda^{2}}\left(\lambda h_{ij}+W_{i}W_{j}\right)\quad{\rm and}\quad b_{i}=-\frac{W_{i}}{\lambda},$ (2)
where $W_{i}=h_{ij}W^{j}$ and $\lambda=1-||W||^{2}$. The weak drift condition, which implies $\lambda>0$, is essential to this derivation. It is worth noticing [9] that
$||b||^{2}_{F}=a^{ij}b_{i}b_{j}=||W||^{2},$ (3)
where $a^{ij}$ is the inverse of the metric tensor $a_{ij}$. Hence, the weak drift also implies $||b||^{2}_{F}<1$, which is commonly known as the Finsler condition in the Physics literature. As it will become clear in the next section, describing the Finslerian structure of black holes requires extending the Zermelo problem to the critical $(||W||=1)$ and strong $(||W||>1)$ drift cases. Fortunately, both problems have been thoroughly investigated and, for the purposes of this work, conclusively solved in the recent mathematical literature, see [7] for a recent review on this subject.
The Zermelo problem for a critical drift was solved in [21, 22] through the analysis of geodesics in the singular limit $\lambda=1$ of the Randers metric (1). The resulting metric
$F(\mathbf{x},d\mathbf{x})=\frac{h_{ij}dx^{i}dx^{j}}{2W_{j}dx^{j}},$ (4)
defined on the half-tangent space
$A=\{d\mathbf{x}\in T_{\mathbf{x}}\mathcal{M}\ :\ W_{j}dx^{j}>0\},$ (5)
is referred to as the Kropina metric in the mathematical literature. Here, the germ of the one-directional membrane behavior which characterizes horizons becomes apparent: solutions (minimal time trajectories) exist only when their tangent vectors $d\mathbf{x}=(dx^{i})$ are positively aligned with the drift $W^{j}$.
The strong drift case, on the other hand, was solved in [23], where it is shown that minimal time trajectories correspond to the geodesics of the so-called Lorentz-Finsler metrics
$F(\mathbf{x},d\mathbf{x})=-\frac{1}{\lambda}\left(W_{i}dx^{i}\mp\sqrt{\left(W_{j}dx^{j}\right)^{2}+\lambda h_{ij}dx^{i}dx^{j}}\right),$ (6)
defined on the conic region of the tangent space $T_{\mathbf{x}}\mathcal{M}$ given by
[…]
$\left.\lambda h_{ij}dx^{i}dx^{j}+\left(W_{j}dx^{j}\right)^{2}>0,\ W_{j}dx^{j}>0\right\}.$
We shall refer to the minus and plus cases, respectively, as first and second Lorentz-Finsler metrics. The strong drift condition $\lambda<0$ guarantees that both metrics in (6) are positive definite. Additionally, as in the critical case, the problem exhibits a one-directional behavior: a solution exists only if the trajectories are positively aligned with the drift, and moreover their instantaneous velocities lie within the conic region (7).
The unit ball on $T_{\mathbf{x}}\mathcal{M}$ defined from the condition $F(\mathbf{x},d\mathbf{x})=1$, called the indicatrix of the Finsler metric, encodes important information on the geodesics. In particular, the indicatrix determines the possible directions $d\mathbf{x}\in T_{\mathbf{x}}\mathcal{M}$ of geodesics starting at the point $\mathbf{x}\in\mathcal{M}$. In our case, for all metrics, including the second Lorentz-Finsler, the indicatrices are given by
$h_{ij}(dx^{i}-W^{i})(dx^{j}-W^{j})=1.$ (8)
In other words, the indicatrices of the Finsler metrics associated with the Zermelo navigation problem on a Riemannian manifold $(\mathcal{M},h_{ij})$ with drift $W^{i}$, regardless of the value of $||W||$, correspond to the unit balls of the Riemannian metric $h_{ij}$, with their centers shifted by the vector $W^{i}$. Randers had already recognized this type of eccentricity in his pioneering work [4]. This result will be essential for the analysis of the causal structure of black holes in Section III.
II.1 Geodesic matching
The mathematical description of the causal structure of Finsler spaces is both comprehensive and sophisticated, see, e.g.[7, 26, 27, 28, 29]. For our purposes here, we need only consider the matching of geodesics across the three cases of Zermelo’s problems. Differentiability issues will not be discussed, as all functions are assumed to be sufficiently smooth. Let $W^{i}$ be a generic smooth vector field drift on a Riemannian manifold $(\mathcal{M},h_{ij})$ and assume there are regions of $\mathcal{M}$ characterized by weak ($||W||<1$) and strong $(||W||>1)$ drifts. The boundary between these regions is a co-dimension 1 submanifold $\mathcal{H}$ of $\mathcal{M}$, corresponding to the critical drift case $||W||=1$. The key observation is that geodesics can cross smoothly from the weak to the strong drift regions, provided that the half-space (5) and conic (7) constraints are satisfied in the pertinent regions. This is possible because the Randers (1), Kropina (4), and first Lorentz-Finsler (6) metrics all give rise to the same Finslerian Lagrangian action $S=\int F(\mathbf{x},d\mathbf{x})$ for geodesics. Notice that the half-space and conic constraints are essential to assure that we have positive Lagrangians $F(\mathbf{x},d\mathbf{x})$ and consequently well-defined geodesic problems, on $\mathcal{H}$ and in the strong drift region.
A concrete example is instructive here. Although simple, the following scenario is paradigmatic, since it corresponds to the river model of the Schwarzschild black hole [25], which we discuss in the next section. Let us examine the three-dimensional Euclidean space $(\mathbb{R}^{3},\delta_{ij})$ with standard Cartesian coordinates $(x,y,z)$, endowed with an infalling radial drift such that $W_{i}dx^{i}=-\sqrt{2M/r}\,dr$, where $r^{2}=x^{2}+y^{2}+z^{2}$. In this case, $\mathcal{H}$ is a spherical surface of radius $r=2M$, identified with the event horizon of the black hole. Its interior and exterior correspond, respectively, to the strong and weak drift regions.
Given the spherical symmetry of the problem, we can consider, without loss of generality, the Randers (1), Kropina (4), and the first Finsler-Lorentz (6) metrics along the $x$ axis. A unified description is given by
$F(\mathbf{x},d\mathbf{x})=\frac{1}{\lambda}\left(\sqrt{dx^{2}+\lambda(dy^{2}+dz^{2})}+\sqrt{\frac{2M}{r}}dx\right),$ (9)
with $\lambda=1-2M/r$. The three metrics correspond, respectively, to $r>2M$ ($\lambda>0$), $r=2M$ ($\lambda=0$), and $r<2M$ ($\lambda<0$). Hence, a unique Finslerian Lagrangian is defined over the entire range of $r$, which implies that geodesics can smoothly cross the submanifold $r=2M$ provided the half-space (5) and conic (7) constraints are satisfied in the respective regions.
The Finslerian indicatrix (8) associated with (9) reduces to the simple expression
$\left(dx+\sqrt{\frac{2M}{r}}\right)^{2}+dy^{2}+dz^{2}=1,$ (10)
which can be readily identified as the unit sphere with a displaced center. Notice that (10) is valid for all metrics, including the second Lorentz-Finsler, and we have
$-1-\sqrt{\frac{2M}{r}}\leq dx\leq 1-\sqrt{\frac{2M}{r}},$ (11)
for all cases. In the exterior (weak drift) region, characterized by $\sqrt{2M/r}<1$, both ingoing ($dx<0$) and outgoing ($dx>0$) geodesics are allowed. For $r\to\infty$, the indicatrix sphere is centered at the origin, implying that there is no preferred direction, as one would expect in asymptotically flat cases. On the other hand, on the critical submanifold $r=2M$, there are no outgoing geodesics, as $-2\leq dx\leq 0$. Finally, for the interior (strong drift) region, where $\sqrt{2M/r}>1$, all geodesics are ingoing since $dx<0$. Fig. 1 illustrates the three cases.
$x$$r=2M$(event horizon)
Figure 1: Depiction of the Finslerian indicatrices for the three drift regimes of the Schwarzschild case along the $x$ axis in the equatorial plane. In all cases, the indicatrices are unit spheres, with their centers displaced in the $x$ direction according to (10). Right: The weak drift regime $r>2M$, in which null geodesics starting from a given point can propagate in all directions. Center: The critical regime $r=2M$, in which the propagation of null geodesics is restricted to the directions that are positively aligned with the drift according to the half-space restriction (5). In this case, the half-space is tangent to the critical submanifold $\mathcal{H}$ (the event horizon), implying that there is no outgoing null geodesics starting on $\mathcal{H}$. Left: The strong drift regime. Here, propagation is similarly limited to directions positively aligned with the drift and, moreover, constrained to lie within the conic region (blue dotted lines) according to (7). Notice that as the starting point of the geodesics approaches the singularity at $r=0$, the displacement of the center of the unit sphere increases, resulting in more acute cones. In the vicinity of the central singularity, all geodesics are rapidly dragged towards it, and there is no way to prevent them from reaching $r=0$, in perfect agreement with the understanding that $r=0$ is a future timelike singularity.
The resemblance of this figure to the river model of black holes [25] is evident. We emphasize that all information regarding geodesics, including their directional dependence, is encoded in the Finsler metric (9), which holds for all $r>0$.
III Black holes
We now proceed to examine the causal structure of black holes through the Finslerian framework. Our first remark is that the forward-in-time ($dt>0$) null geodesics ($ds=0$) of any stationary spacetime whose line element is cast in the Painlevé-Gullstrand form [30]
$ds^{2}$ $=$ $g_{\mu\nu}dx^{\mu}dx^{\nu}$
$=$ $-dt^{2}+h_{ij}\left(dx^{i}-W^{i}dt\right)\left(dx^{j}-W^{j}dt\right),$
with both time-independent spatial metric $h_{ij}$ and drift $W^{i}$, are determined by a positive definite Finsler metric $dt=F(\mathbf{x},d\mathbf{x})$ of the Randers type characterized by (1) and (2). In the standard river model of black holes [25], $h_{ij}$ is assumed to be an Euclidean metric, and in this case the metric (III) is said to be of strong Painlevé–Gullstrand form [31]. We note, however, that this restriction on $h_{ij}$ is not required in our analysis. The metric (III), up to a conformal term, is referred to as the Zermelo form of a stationary spacetime in the recent Physics literature [9, 11] . We also remark that the spacetimes for which the main recent mathematical results were established [7], possessing a space-transverse Killing vector field, have metrics precisely of the form (III), with $g_{00}=1-||W||^{2}$ (note that $g_{00}$ in our notation corresponds to $\Lambda$ in the notation of [7]). Building on the discussion of Zermelo’s problem in Section II, we can describe the propagation of null geodesics of (III) using the underlying Finsler metric $dt=F(\mathbf{x},d\mathbf{x})$, irrespective of the value of $\lambda=1-h_{ij}W^{i}W^{j}$. The value of $\lambda$, in turn, determines whether the Finsler metric takes the form of Randers (1), Kropina (4), or Lorentz-Finsler (6). Accordingly, it also identifies the relevant Zermelo problem associated with the black hole spacetime under investigation. We will now consider static and stationary black holes separately.
III.1 Static black holes
The generic line element for a static spherically symmetrical spacetime reads
$ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}=-f(r)d\tau^{2}+\frac{dr^{2}}{f(r)}+R^{2}(r)dS^{2},$ (13)
where $dS^{2}$ stands for the usual metric on the unit sphere. Asymptotic flatness requires $R\sim r$ and $f\sim 1$ for sufficiently large $r$. The zeros of $f(r)$ correspond to Killing horizons of this spacetime, with the outermost zero corresponding to the event horizon of the black hole. We can cast (13) in a Painlevé-Gullstrand form by introducing the new time variable [30]
$dt=d\tau+\frac{1}{f}\sqrt{1-f}dr,$ (14)
leading to a metric of the form (III) with drift given by
$W_{i}dx^{i}=-\sqrt{1-f}dr$ (15)
and spatial Riemannian metric
$d\ell^{2}=h_{ij}dx^{i}dx^{j}=dr^{2}+R^{2}(r)dS^{2}.$ (16)
Note that, unless $R(r)=r$, the associated Zermelo navigation problem is defined on a curved manifold. Additionally, we have $\lambda=f$ and the underlying Finsler metric reads
$dt\!=\!F(x,dx)\!=\!\frac{1}{f}\left(\!\sqrt{1-f}dr+\sqrt{dr^{2}+fR^{2}dS^{2}}\right)\!.\!$ (17)
The importance of the Painlevé-Gullstrand coordinates becomes evident now. They can be smoothly extended across the horizon ($f=0$), yielding the Finsler metric (17), which remains well defined throughout the entire domain of the radial coordinate $r$ where the Painlevé-Gullstrand metric is also defined. Consequently, the Finsler metric (17) is also well defined for all values of $\lambda$. Besides, in the strong drift case $(\lambda<0)$, we also have the second Lorentz-Finsler metric. We now turn to some explicit examples for illustration.
The Schwarzschild case analyzed in the previous section can be generalized to other black holes described by (13). In fact, any black hole with metric such that $R(r)=r$ will correspond to a Zermelo problem defined on a flat tridimensional manifold. The corresponding indicatrices, determined by (8), are also given by spheres with displaced center, such that
$\left(dx+\sqrt{1-f}\right)^{2}+dy^{2}+dz^{2}=1,$ (18)
which implies
$-1-\sqrt{1-f}\leq dx\leq 1-\sqrt{1-f}.$ (19)
As in the Schwarzschild case, limiting the analysis to the $x$ axis does not result in any loss of generality, owing to the spherical symmetry of the problem.
Many other physically relevant spacetimes belong to this class of Euclidean Zermelo problems. For instance, the Rindler metric considered in [11], characterized by the constant radial acceleration $\gamma$, corresponds to $f=1-2\gamma r$. For the Reissner-Nordström black hole, characterized by the electric charge $Q$ such that $|Q|<M$, we have
$f(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}=\frac{(r-r_{-})(r-r_{+})}{r^{2}},$ (20)
where $r_{\pm}=M\pm\sqrt{M^{2}-Q^{2}}$. The Reissner-Nordström-de Sitter case can also be accommodated here, see [11] for additional examples and further details. We emphasize that there are static black holes, such as dilaton black holes [33], which necessarily require a curved spatial manifold. The only difference with respect to the flat cases is that in these non-Euclidean scenarios, the indicatrix (8) will be an eccentric ellipsoid instead of an eccentric sphere as in (18). Still, ultimately, the qualitative results will remain the same.
It is worth examining the Reissner-Nordström black hole in greater detail. It is clear from (20) that we have two Killing horizons, but there is also another distinguished radius, namely $r=r_{0}=Q^{2}/2M$. The Painlevé-Gullstrand metric for the Reissner-Nordström black hole is defined only for $r\geq r_{0}$, since $f(r)>1$ for $r<r_{0}$ and the square root in (14) fails to be real.
Given that $r_{0}$ is half of the harmonic mean of the two horizons, it follows that $r_{0}<r_{-}$. Hence, the Painlevé-Gullstrand coordinates for the Reissner-Nordström black hole effectively extend across its inner horizon $r=r_{-}$. For $r>r_{+}$, $\lambda=f(r)>0$ and we are in the weak drift regime of the Zermelo problem. Similarly to the exterior region of the Schwarzschild black hole of the previous section, in this region we can have both ingoing ($dx<0$) and outgoing ($dx>0$) null geodesics.
For $r=r_{+}$ and $r=r_{-}$, we have $f(r)=0$, which correspond to the critical Zermelo case. The associated geodesics are such that $2\leq dx\leq 0$, as in the case of the Schwarzschild event horizon. In the region between the two Killing horizons, $r_{-}<r<r_{+}$, where $f(r)<0$, we have the strong drift case. This regime is similar to the one occurring in the interior region of the Schwarzschild black hole, where we have only ingoing geodesics since $dx<0$. The novelty for the Reissner-Nordström black hole is that, for $r_{0}\leq r<r_{-}$, we have again the weak drift regime corresponding to $f(r)>0$. In this regime, both ingoing ($dx<0$) and outgoing ($dx>0$) null geodesics are allowed, but they remain entirely confined within the inner horizon. This indicates that null geodesics starting within the inner (Cauchy) horizon of the Reissner-Nordström black hole can escape from reaching the spacelike singularity at $r=0$. This characteristic is well-known in the theoretical study of charged black holes. However, it is generally not considered physically relevant because the electric charge in astrophysical black holes is expected to neutralize quickly [34, 35]. Additionally, there are inherent instabilities associated with the Cauchy horizon, see [36, 37, 38].
We remark that the Painlevé-Gullstrand coordinates for the Reissner-Nordström black hole are defined only for $r>r_{0}$. Consequently, the Finslerian description of its causal structure is limited by the same restriction. However, this does not pose a major concern for our purposes, as it is related to the presence of a negative local mass, which effectively prevents ingoing geodesics from reaching the central singularity at $r=0$, a well-known issue with the Painlevé-Gullstrand coordinates. For further details, see [39]. For an interpretation of this phenomenon in the context of the river model of black holes, refer to [25].
III.2 Stationary black holes
The case of stationary black holes is significantly richer and physically more interesting. Here, we focus exclusively on the Kerr metric, which describes electrically neutral rotating black holes, thus avoiding unnecessary complications such as the negative local mass of Reissner-Nordström black holes discussed above.
It is well known that the Kerr metric cannot be cast in a strong Painlevé–Gullstrand form [31]. In fact, its representation within the river model paradigm relies on the Doran coordinates [32], requiring the introduction of a “twist” drift. In this work, we consider the Painlevé–Gullstrand form of the Kerr metric proposed by Natario in [40], which corresponds to the Zermelo metric (III) with the Riemannian spatial metric
$h_{ij}dx^{i}dx^{j}=\frac{\rho^{2}}{\Sigma}dr^{2}+\rho^{2}d\theta^{2}+\Sigma\sin^{2}\theta\left(d\phi+\delta d\theta\right)^{2},$ (21)
and drift $W^{i}$ such that
$W_{i}dx^{i}=\frac{\rho^{2}}{\Sigma}vdr+\Sigma\sin^{2}\theta\Omega d\phi,$ (22)
with
$v=-\frac{\sqrt{2Mr(r^{2}+a^{2})}}{\rho^{2}},\quad\Omega=\frac{2Mra}{\rho^{2}\Sigma},$ (23)
$\rho^{2}=r^{2}+a^{2}\cos^{2}\theta,\quad\Sigma=r^{2}+a^{2}+\frac{2Mra^{2}}{\rho^{2}}\sin^{2}\theta,$ (24)
and
[…]
We remark that the Kerr metric describes a black hole, spinning with angular velocity $\Omega$, only if $|a|\leq M$; otherwise it corresponds to a naked singularity.
Notice that (21) and (22) are well defined for all positive $r$. Natario’s coordinates $(t,r,\theta,\phi)$ differ from the usual Boyer-Lindquist in the time variable, which is not relevant to our analysis, and in the azimuthal angle. The usual Boyer-Lindquist azimuthal coordinate $\phi_{BL}$ is given by (see [40] for details)
$d\phi_{BL}=d\phi+\frac{\rho^{2}}{\Delta}v\Omega dr+\delta d\theta,$ (26)
where $\Delta=r^{2}-2Mr+a^{2}$. Since the coordinates $r$ and $\theta$ are the same in Boyer-Lindquist and Natario’s coordinates, together with the fact (21) and (22) do not depend on $t$ and $\phi$ (as expected due to axisymmetry and stationarity), the ergosurfaces and the horizons of the Kerr metric are described by the same equations in both coordinate systems. In fact, the outer and inner ergosurfaces are given by the solutions of $\rho^{2}=2Mr$, whereas the outer and inner event horizons are given by the roots of $\Delta=0$.
The Finsler indicatrix for (21) and (22) reads
$\frac{\rho^{2}}{\Sigma}\left(dr-v\right)^{2}+\rho^{2}d\theta^{2}+\Sigma\sin^{2}\theta(d\phi+\delta d\theta-\Omega)^{2}=1.$ (27)
Geometrically, it corresponds to an ellipsoid with displaced center in the tangent space of the spatial sections of (III) such that
$-\frac{\sqrt{\Sigma}}{\rho}\left(1-\frac{\rho v}{\sqrt{\Sigma}}\right)\leq dr\leq\frac{\sqrt{\Sigma}}{\rho}\left(1+\frac{\rho v}{\sqrt{\Sigma}}\right).$ (28)
Given the relation
$\frac{\rho v}{\sqrt{\Sigma}}=-\sqrt{1-\frac{\Delta}{\Sigma}},$ (29)
we deduce that $dr\leq 0$ when $\Delta=0$, which implies that no geodesic can be outgoing at either horizon. Between the horizons, the equations above imply $dr<0$, meaning all geodesics are ingoing. Hence, the qualitative behavior of the event horizons of a Kerr black hole is analogous to what was observed previously for static black holes. Additionally, similarly to the Reissner-Nordström case, null geodesics can avoid the ring singularity at $r=0$ when $0\leq r\leq r_{-}$, although they remain confined to the interior of the inner horizon, with the same caveats concerning the physical relevance of this picture due to instabilities issues of the inner horizon [38].
The analysis of the ergosurfaces is more involved. Consider a geodesic which is locally aligned with the drift vector $W^{i}$, characterized by
$dr=\alpha v,\quad d\theta=0,\quad d\phi=\alpha\Omega.$ (30)
In this case, the equation (27) that determines the indicatrix simplifies to
$(\alpha-1)^{2}||W||^{2}=1,$ (31)
where
$||W||^{2}=\frac{\rho^{2}v^{2}}{\Sigma}+\Sigma\sin^{2}\theta\Omega^{2}=\frac{2Mr}{\rho^{2}}.$ (32)
From this relation, we see that the ergosurfaces correspond effectively to the critical drift case $||W||=1$. Since the solutions of (31) are $\alpha=1\pm||W||^{-1}$, we have that on the ergosurfaces we effectively always have a trapped direction, which coincides with the local direction of $W^{i}$. More precisely, the trapped propagation direction ($dx^{i}=0$) for null geodesics corresponds to the origin of $T_{\mathbf{x}}\mathcal{M}$, and the origin belongs to the indicatrix only in the critical case. However, unlike the static case, for rotating black hole this trapped direction is not the radial one.
Fig. 2 illustrates the results above for geodesics confined to the equatorial plane $\theta=\pi/2$ and starting along the $x$ axis. In this case, the indicatrix equation (27) reads
$\frac{r^{2}+a^{2}}{\Sigma}\left(dx+\sqrt{\frac{2M}{r}}\right)^{2}$
$+\frac{\Sigma}{r^{2}+a^{2}}\left(dy-\frac{2Ma\sqrt{r^{2}+a^{2}}}{r\Sigma}\right)^{2}=1,$ (33)
where we have used explicitly the fact that Natario’s coordinates are also of the oblate spheroidal type and, hence, on the equatorial plane the relations $r^{2}+a^{2}=x^{2}+y^{2}$ and $\tan\phi=y/x$ hold.
$x$$r=2M$(outer ergosurface)$r=M+\sqrt{M^{2}-a^{2}}$(outer horizon)
Figure 2: Depiction of the Finslerian indicatrices for the ergosphere and the outer event horizon along the $x$ axis in the equatorial plane of a Kerr black hole ($a>0$) in Natario’s coordinates (21). When $a<0$, the figure is reflected with respect to the $x$ axis, corresponding to the transformation $y\to-y$ in (III.2). In all cases, the indicatrices are horizontal ellipses, with their centers displaced according to (III.2). Notice that the centers of the indicatrices are also displaced in the azimuthal direction (vertical in the figure) due to the well-known frame-dragging effect. Right: The weak drift regime, in which null geodesics starting from a given point can propagate in all directions. Center: The critical regime corresponding to the ergosurface. The propagation of null geodesics is restricted to directions that are positively aligned with the drift, as shown in equation (5). In contrast with the Schwarzschild ($a=0$) case, the half-space of possible propagation directions (bounded by the blue dotted line) is not tangent to the ergosphere, meaning that null geodesics starting on the ergosphere can still escape to infinity. Nevertheless, they are subjected to $d\phi_{BL}\geq 0$, meaning that null geodesics are corotating with the black-hole in the usual Boyer-Lindquist coordinates. This is the only case where we can have trapped directions $(dx^{i}=0)$. Left: The event horizon. In this case, the propagation is similarly limited to the directions that are positively aligned with the drift and, moreover, constrained to lie within the conic region outlined in (7) (blue dotted line). Notice the cone is such that $dr\leq 0$ on the event horizon, meaning that no null geodesics can escape from the interior of the horizon.
The solutions of the equation above, corresponding to the indicatrices for rotating black holes, are “horizontal” ellipses with displaced center, analogously to the static case of the previous section. Moreover, the closer the starting point of the null geodesics is to the ring singularity at $r=0$, the more elongated the ellipse becomes in the $x$ direction. Furthermore, in contrast with the previous static cases, there is now also a displacement in the vertical direction, which is clearly related to the well-known frame-dragging effects.
It is convenient to consider the azimuthal Boyer-Lindquist angle (26) to understand the distinctive properties of the ergosurfaces and their interior region. For the trapped geodesic described by (30), we have
$d\phi_{BL}=\alpha\Omega\left(1+\frac{\rho^{2}v^{2}}{\Delta}\right),$ (34)
demonstrating that the ergosphere corresponds to the frontier of the “corotating” region (ergosphere) where $d\phi_{BL}$ and $\Omega$ have the same sign, i.e. $d\phi_{BL}\geq 0\,(\leq 0)$ for $\Omega>0\,(<0)$. Notice (34) explicitly shows the well-known fact that the azimuthal Boyer-Lindquist angle is well defined only in the exterior region of the event horizon, in clear contrast with Natario’s azimuthal coordinate.
Finally, it is worth noticing that in the limit of small angular momentum, where only the linear terms in $a$ are considered, the spatial metric (21) simplifies to a flat metric. In this case, we recover the Zermelo navigation problem associated with the slowly rotating Kerr metric discussed in reference [11]. The Finslerian indicatrices in this scenario are qualitatively similar to those shown in Fig. 2, but with the event horizon and the ergosurface positioned arbitrarily close to each other.
IV Final remarks
We have shown, through explicit and relevant physical examples, that recent mathematical findings (see [7] for a recent review) regarding the Zermelo navigation problem and its underlying Finsler geometry can provide a complete description of the causal structure of black holes, answering affirmatively the question posed in Ref. [11] about the extension of the Finsler structure to the interior of black holes. Specifically, the Kropina metric, associated with the critical Zermelo navigation problem, can fully characterize the causal properties of ergosurfaces in stationary black holes and event horizons in the spherically symmetric case. On the other hand, the Lorentz-Finsler metric, which applies to the case of a strong Zermelo problem, describes the causal structure of the interior (non-stationary) region of black holes. This complete Finslerian description offers an enhanced visual understanding of the river model of black holes, presenting the behavior of null geodesics on the ergosurfaces, horizons, and their interior regions in novel geometric terms. There are many other recent mathematical results that may have physical applications within this framework; these topics are currently under investigation.
[…]
Our second and final remark concerns the analog gravity program, a robust framework that explores the emergence of effective curved spacetime geometries in non-gravitational systems [41, 42, 43]. The study of Finsler geometries can be further advanced by investigating their connections to analog spacetimes, such as those realized in hydrodynamic flows [9, 20, 44, 45], Bose-Einstein condensates [44, 45, 46], and birefringent crystals [15, 47]. As we have argued, the interplay between Spacetime Physics and Finsler geometries becomes particularly evident when the metric is recast in the Painlevé-Gullstrand form. The Painlevé-Gullstrand form not only emphasizes the causal structure of the underlying spacetime, but also naturally aligns with the Zermelo framework. In analog spacetimes, this alignment is not incidental since the Painlevé-Gullstrand form is particularly suitable for systems with a preferred frame, such as fluids or optical media. By establishing a direct connection between Finsler geometries and the causal structure of black hole interiors through explicit constructions, our work lays the foundation for deeper investigations into analog horizons and ergosurfaces, offering fresh perspectives on Finslerian concepts within the framework of analog gravity.
Acknowledgments
This study was financed in part by the São Paulo Research Foundation (FAPESP, Brazil) - Process Numbers 2022/15371-3 (HR), 2022/08335-0 (MR), 2021/09293-7 (AS). The authors also acknowledge partial support from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq, Brazil), Grants 315991/2023-2 (MR) and 306785/2022-6 (AS). AS also wishes to thank Vitor Cardoso and José S. Lemos for the warm hospitality at the Center for Astrophysics and Gravitation of the University of Lisbon, where part of this work was done.